OpenStudy (anonymous):

can i get an example for a problem of solving an absolute value equation, please?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

the absolute value of x is 6. so x could be either 6 or -6

OpenStudy (anonymous):

\[|x-3|=5\] \[x-3=5\] \[x-3=-5\] \[x-3=5 , x=8\] \[x-3=-5, x=-2\]

OpenStudy (anonymous):

there are two. do you have one you would like to do?

OpenStudy (anonymous):

yeah, the |x-3| = 5 one.

OpenStudy (anonymous):

the problem is, I don't quite remember how to do the process. Could one help.

OpenStudy (anonymous):

just remember when you remove the absolute value signs you have to write two separate equations

OpenStudy (anonymous):

this is clear yes? because if i know \[|x|=78\] then x could be 78 but it could also be -78

OpenStudy (anonymous):

yeah, I just looked that up, and it made sense. Now what if there's two variables, like |x-3| = 5, would the answer be, "2 or -2" ?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

what? why?

OpenStudy (anonymous):

because if i know |x-3|=5 then what i know is that either x-3=5 or x-3=-5

OpenStudy (anonymous):

the solution to the first one is x = 8, not x = 2. the solution to the second one IS x=-2

OpenStudy (anonymous):

Ok. I kind of see what you did... Could you give me another one amongst those lines, so I can try it?

OpenStudy (anonymous):

sure try |x+3|=7

OpenStudy (anonymous):

The answer to the first one should be 4, and the second one, being (|x+3| = -7) would be -10?

OpenStudy (anonymous):

yes . got it. write to equations, solve each individually

OpenStudy (anonymous):

i meant write TWO equations

OpenStudy (anonymous):

Yeah, for sure! (: It's just I'm a bit slow at first, but once I get it, I beast through all of them, lol. Btw, are there any equations like that, with THREE variables?

OpenStudy (anonymous):

not usually. do you have one?

OpenStudy (anonymous):

No, I just wanted to know so I could be prepared in case I get one of those.

OpenStudy (anonymous):

you could have something like this: |x+2|=2x-5

OpenStudy (anonymous):

and how would you solve that?

OpenStudy (anonymous):

in which case you still have to write two equations but you have to be very very careful

OpenStudy (anonymous):

you would write two equations. one would be: x+2=2x-5

OpenStudy (anonymous):

the other would be: x+2=-(2x-5)

OpenStudy (anonymous):

which of course means x+2=-2x+5

OpenStudy (anonymous):

be careful when you take the negative to remember to put it in parentheses

OpenStudy (anonymous):

alright, and so from there, what step is next?

OpenStudy (anonymous):

just solve like any linear equation. \[x+2=2x-5\] \[2+5=2x-x\] \[7=x\]

OpenStudy (anonymous):

or \[x+2=-(2x-5)\] \[x+2=-2x+5\] \[x+2x=5-2\] \[3x=3\] \[x=1\]

OpenStudy (anonymous):

OH, OK!

OpenStudy (anonymous):

hey do you have time, ? because i have another question...

OpenStudy (anonymous):

and btw, thanks a lot.

OpenStudy (anonymous):

AWH, no satellite, dont forsake me right now! I got aaaaaa big TEST tomorrow!

OpenStudy (anonymous):

one more and then i turn into a pumpkin

OpenStudy (anonymous):

what's that mean?

OpenStudy (anonymous):

go ahead and ask

OpenStudy (anonymous):

never saw or read cinderella?

OpenStudy (anonymous):

oh, yeah, aha.... but its only 10:44, doesn't that happen at 12?

OpenStudy (anonymous):

eastern standard time

OpenStudy (anonymous):

Interesting. Ok, anyways, sorry for delaying your transformation, but could you give me some heads up on solving prequalities?

OpenStudy (anonymous):

?? ptequalities?

OpenStudy (anonymous):

even spell check doesn't like that

OpenStudy (anonymous):

I don't know, the weird lady put one of the subjects we need to study, under "Solving prequaltities and graphing solution." Any ideas?

OpenStudy (anonymous):

sorry never heard of them.

OpenStudy (anonymous):

google hasn't either so i don't feel bad

OpenStudy (anonymous):

Yeah, it doesn't even show up on google. !

OpenStudy (anonymous):

Alright, well, you've been of great help! And I appreciate it. Thank you, who ever you are, and have a good one!

OpenStudy (anonymous):

gnight.